A random sample is obtained from a population with μ = 120 and σ = 20, and a treatment is administered to the sample. Which of the following outcomes would be considered noticeably different from a typical sample that did not receive the treatment?

a. n = 36 with M = 121
b. n = 36 with M = 123
c. n = 144 with M = 121
d. n = 144 with M = 124
Answer #1

Solution:

          we are given that : a random sample is obtained from a population with μ = 120 and σ = 20, and a treatment is administered to the sample.

         We have to find from the following outcomes that would be considered noticeably different from a typical sample that did not receive the treatment.

a. n = 36 with M = 121
b. n = 36 with M = 123
c. n = 144 with M = 121
d. n = 144 with M = 124

Thus we need to find z score for each :

z score formula :

z = \frac{M-\mu}{\sigma /\sqrt{n}}

Part a)

n = 36 with M = 121

z = \frac{M-\mu}{\sigma /\sqrt{n}}

z=\frac{121-120}{20/\sqrt{36}}

z=\frac{1}{20/6}

z=\frac{1}{3.333333}

z=0.30

b) n = 36 and M = 123

z=\frac{123-120}{20/\sqrt{36}}

z=\frac{3}{20/6}

z=\frac{3}{3.333333}

z=0.90

c) n = 144 and M = 121

z=\frac{121-120}{20/\sqrt{144}}

z=\frac{1}{20/12}

z=\frac{1}{1.666667}

z=0.60

d)   n = 144 and M = 124

z=\frac{124-120}{20/\sqrt{144}}

z=\frac{4}{20/12}

z=\frac{4}{1.666667}

z=2.40

Thus z = 2.40 is for n = 144 with M = 124,

thus outcome d. n = 144 with M = 124 would be considered noticeably different from a typical sample that did not receive the treatment.